# https://xiaojiu-z.github.io/Stable-Makeup.github.io/

### v1

@@ G Further Exploration of Data Construction
Currently, diffusion-based methods have demonstrated remarkable performance in terms of fidelity and diversity, making them a crucial tool for achieving high-quality and diverse data in step 2 of our data construction process. In addition to the inversion-based image editing methods we employed, we could also choose to use existing identity customization methods, such as InstantID[44], which do not require test-time fine-tuning. However, as shown in figure 7, when using these diffusion-based methods for editing facial images, visible changes in facial structure may occur during makeup generation guided by text. For example, the mouth shape of these generated images cannot be exactly the same. Given the current limitations of diffusion-based methods in terms of fine-grained control, our proposed content and structure decoupling strategy is particularly important for maintaining the facial expression and structure.
Figure 7: Visual results of different data construct pipeline.
H Designing Choice of Detail-Preserving Makeup Encoder
@@
In this section, we conducted a comparative study between different mapping structure setting under the same training conditions to demonstrate that mapping without preserving spatial information struggling to align and preserving makeup details at a fine-grained level, and to validate the effectiveness of our proposed method. We first introduced our experimental settings in Section 8. Subsequently, we conduct visual comparison in Sec. H.2 to demonstrate the pivotal role of our proposed method in preserving and aligning the makeup details. Additionally, in Sec. H.3 we further demonstrated the effectiveness of our proposed method in preserving and aligning the makeup details by visualizing the attention maps of the makeup cross-attention layers in U-Net.
H.1 Experimental Setup
@@
Figure 8: The structural composition of the five distinct mapping schemes. We omitted the drawing of the global vector (cls token from CLIP) for the sake of brevity and clarity.
@@
  • Self-Attention mapping: We utilized our proposed self-attention mapping in section 3.1.2. The projected dimension is the same as the original dimension (1, 257$\times$12, 1024).
@@
Figure 9: Visual comparison results under the same training step but with different mapping settings.
H.2 Comparison
@@
H.3 Visualization of Attention maps
Figure 10: Visualization of attention maps with different mapping settings at the 50,000th training step.
In order to conduct a more comprehensive comparison between spatial linear mapping and self-attention mapping, we visualized the attention maps of the makeup cross-attention layers under both configurations at the 50,000th training step in this experiment. Specifically, we extract and aggregate the attention matrices calculated from the patch features of the source facial region (Query) and the multi-scale makeup embeddings (Key-Value), based on the index of the facial position, and visualize them at different layers, as shown in Figure 10 (a). From the attention maps, we can see that the self-attention layer can effectively weight the features, resulting in high response for the facial patch features in between the source image and the makeup image, while low response between the other positions in the makeup image. However, when we replace the self-attention mapping with spatial linear mapping, the facial region of the source image also exhibit high response to the background features, as shown in the attention maps of layer1-4 in Figure 10 (b). These findings highlight the advantages of self-attention mapping in achieving more precise and effective feature weighting, which is crucial for achieving high-quality makeup transfer results.
I More Results
@@
J Broader Impact
The broader impact of diffusion-based makeup transfer methods is significant, as they have the potential to revolutionize the beauty industry by enabling more efficient and personalized makeup application. However, it is important to consider the ethical implications of such technology, including issues related to privacy, consent, and the potential perpetuation of beauty standards. We do not condone the unauthorized use of our method to edit the makeup of others’ photos. As with any emerging technology, we urge caution in the use of diffusion-based makeup transfer methods and further consideration of their ethical and legal implications.
K Limitations
@@
Figure 11: More results of makeup transfer for realistic faces and cross-domain faces.
Figure 12: More results of video makeup transfer and makeup guided text-to-image generation.

### v2

@@ Appendix D Further Exploration of Data Construction
Currently, diffusion-based methods have demonstrated remarkable performance in terms of fidelity and diversity, making them a crucial tool for achieving high-quality and diverse data in step 2 of our data construction process. In addition to the inversion-based image editing methods we employed, we could also choose to use existing identity customization methods, such as InstantID(Wang et al., 2024), which do not require test-time fine-tuning. However, as shown in figure 7, when using these diffusion-based methods for editing facial images, visible changes in facial structure may occur during makeup generation guided by text. For example, the mouth shape of these generated images cannot be exactly the same. Given the current limitations of diffusion-based methods in terms of fine-grained control, our proposed content and structure decoupling strategy is particularly important for maintaining facial expression and structure.
Figure 7. Visual results of different data construct pipelines.
Appendix E Designing Choice of Detail-Preserving Makeup Encoder
@@
In this section, we conducted a comparative study between different mapping structure settings under the same training conditions to demonstrate that mapping without preserving spatial information struggles to align and preserve makeup details at a fine-grained level, and to validate the effectiveness of our proposed method. We first introduced our experimental settings in Section 8. Subsequently, we conduct a visual comparison in Sec. E.2 to demonstrate the pivotal role of our proposed method in preserving and aligning the makeup details. Additionally, in Sec. E.3 we further demonstrated the effectiveness of our proposed method in preserving and aligning the makeup details by visualizing the attention maps of the makeup cross-attention layers in U-Net.
E.1. Experimental Setup
@@
Figure 8. The structural composition of the five distinct mapping schemes. We omitted the drawing of the global vector (cls token from CLIP) for the sake of brevity and clarity.
@@
  • Self-Attention mapping: We utilized our proposed self-attention mapping in the Methodology section. The projected dimension is the same as the original dimension (1, 257$\times$12, 1024).
@@
Figure 9. Results under extreme viewing angles. (please zoom in to better view).
Figure 10. Visual comparison results under the same training step but with different mapping settings.
E.2. Comparison
@@
E.3. Visualization of Attention maps
Figure 11. Visualization of attention maps with different mapping settings at the 50,000th training step.
In order to conduct a more comprehensive comparison between spatial linear mapping and self-attention mapping, we visualized the attention maps of the makeup cross-attention layers under both configurations at the 50,000th training step in this experiment. Specifically, we extract and aggregate the attention matrices calculated from the patch features of the source facial region (Query) and the multi-scale makeup embeddings (Key-Value), based on the index of the facial position, and visualize them at different layers, as shown in Figure 11 (a). From the attention maps, we can see that the self-attention layer can effectively weigh the features, resulting in a high response for the facial patch features in between the source image and the makeup image, while a low response between the other positions in the makeup image. However, when we replace the self-attention mapping with spatial linear mapping, the facial region of the source image also exhibits a high response to the background features, as shown in the attention maps of layers 1-4 in Figure 11 (b). These findings highlight the advantages of self-attention mapping in achieving more precise and effective feature weighting, which is crucial for achieving high-quality makeup transfer results.
Appendix F More Results
@@
Appendix G User study
Table 3. User study. Metrics that are bold represent methods that rank 1st. Our method achieves the highest average score.
Method BeautyGAN CPM PSGAN EleGANt SCGAN SPMT SSAT Ours
Average Score 3.01 3.13 2.95 3.62 2.90 3.35 2.85 4.58
We conducted a user study to compare our method with other methods perceptually. We randomly select 10 source images and 10 reference images from both the M-Wild dataset and the CPM-real dataset. For each evaluation, each user will see one source image with a reference makeup image and one image generated by each method. 60 evaluators were asked to provide a comprehensive score for each generated image based on its makeup consistency, content and structure consistency, using a scale of 1 (worst) to 5 (best). The results of average scores in Table 3 indicate that our method outperforms the comparison methods. Our method achieves the highest average score, which clearly indicates that our method outperforms the comparison methods.
Appendix H Extreme viewing angles
Our approach is capable of processing source images under extreme viewing angles (in Fig. 9).
Appendix I Broader Impact
The broader impact of diffusion-based makeup transfer methods is significant, as they have the potential to revolutionize the beauty industry by enabling more efficient and personalized makeup applications. However, it is important to consider the ethical implications of such technology, including issues related to privacy, consent, and the potential perpetuation of beauty standards. We do not condone the unauthorized use of our method to edit the makeup of others’ photos. As with any emerging technology, we urge caution in the use of diffusion-based makeup transfer methods and further consideration of their ethical and legal implications.
Appendix J Limitations
@@ Appendix I Broader Impact
Figure 12. More results of makeup transfer for realistic faces and cross-domain faces.
Figure 13. More results of video makeup transfer and makeup guided text-to-image generation.